OpenStudy (anonymous):

A data set consists of the following data points: (2,4), (4,7), (5,12) The line of best fit has the equation y =-1.5+2.5x. What does this equation predict for a value of x=3?

Directrix (directrix):

Take the equation for the line of best fit and substitute 3 for x. Crank out the value of y that goes with that x. That will be the predicted value. y =-1.5+2.5x y = -1.5 + 2.5*3 y = ? Post what you get, okay?

OpenStudy (anonymous):

6?

Directrix (directrix):

That is what I got, 6.

OpenStudy (anonymous):

Cool, thanks!